Consider the recurrence equation defining T(n):

T(n)=1 if n-1;

T(n)=T(n-1)+2^n Otherwise .

Show by induction T(n)=2^(n+1)-1
Answer #1

Given That, T (n)=1, if n=o. T(n) = T(n-1) + an other wise show by induction T (n) = 2 Proof by induction nt 2. Base Case - fstep2 (- assume That it is true for Kel T (K) = 2 is true. now we have to prove it for n=k+1 for n=k+1 Step3 K+| K+/ L.H.S: T